OpenStudy (anonymous):

A triangular prism has a volume of 5 m3. If each dimension of the prism is tripled, what is the new volume of the prism?

OpenStudy (anonymous):

45 m3 15 m3 135 m3 405 m3

OpenStudy (mertsj):

Whoops. Each dimension is tripled 0 V=1/2bh(Height of Prism) Let's triple each dimension: V=1/2(3b)(3h)(3Height of Prism) V=27(1/2(bh)(Height of Prism) So the new volume is 27 times the old volume.

OpenStudy (anonymous):

so 135?

OpenStudy (mertsj):

yep

OpenStudy (anonymous):

ok thanks!

OpenStudy (mertsj):

yw